OpenStudy (anonymous):

Could someone double check my answer? Its just solving for y, but I think I messed up. 1. 3x+6y=12 I got y=2-1/2x 2. y-4=-1/2(x+1) For this one I didn't really know where to start so just point me in a direction and then I'll solve it.

OpenStudy (mathstudent55):

The first one is correct.

OpenStudy (anonymous):

Well for the first one we have a simplified (if you divide both sides by 3) version of 3x+6y=12 to x+2y=4

OpenStudy (mathstudent55):

In the second one, just add 4 to both sides.

OpenStudy (anonymous):

So y=(4-x)/2 y=2-(1/2x)

OpenStudy (anonymous):

@mathstudent55 What do I do next in the second one? I'm confused by what I have to do with the parentheses. (Way, way misspelled that :P)

OpenStudy (anonymous):

@codyw00 try using PEMDAS distribute the -1/2 to the parentheses then see what you can do from there

OpenStudy (anonymous):

@amanda7414 Thanks and I've gotten to y=-0.5x+3.5 So is that in y=mx+b formula now?

OpenStudy (anonymous):

correct

OpenStudy (mathstudent55):

\(y-4=-\dfrac{1}{2}(x+1)\) 1. add 4 to both sides: \(y=-\dfrac{1}{2}(x+1) + 4\) The equation is already solved for y, but that may not look very good, so we can try to combine like terms by first distributing the -1/2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Np

OpenStudy (mathstudent55):

\(y=-\dfrac{1}{2}(x+1) + 4\) 2. distribute the -1/2: \(y = -\dfrac{1}{2} x - \dfrac{1}{2} + 4\) 3. combine like terms on right side:

OpenStudy (mathstudent55):

\(y = -\dfrac{1}{2} x - \dfrac{1}{2} + \dfrac{8}{2}\) \(y = -\dfrac{1}{2}x + \dfrac{7}{2} \)